2. Ifanz converges at z = z0, prove that anz absolutely converges in the disk D(0, |zo|).
Answer #1

Suppose the power series anan converges for some z0 with z0 ̸= 0.

Then its terms converge to zero, so they are bounded and there exists M ≥ 0 such that

|a_nz_0^n|\leq M for n = 0, 1, 2, . . . .

If |z| < |z0|, then

Mr > ا ا = = | luewhere r < 1

Comparing the power series with the convergent geometric series \sum_{n=0}^{\infty }Mr^n , we see that D is absolutely convergent. Thus, if the power series converges for some z0 , then it converges absolutely for every z with |z| < |z0|.

So, it absolutely converges in the disk D(0,|z0|).